3x²y²dx + 4(x³y -3) dy = 0
1
Expert's answer
2020-10-14T13:37:41-0400

"3x^2y^2dx+4(x^3y-3)dy=0"

integrating factor is "y^2:"

"3x^2y^4dx+4x^3y^3dy-12y^2dy=0"

The left part is the full differential of the next function:

"x^3y^4-4y^3=C"